Circular No. 5642

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bf IN NGC 930 Jean Mueller reports her discovery, on a blue plate taken Oct. 2 by C.

Brewer, J. D. Mendenhall, and herself with the Oschin Schmidt Telescope in the course of the second Palomar Sky Survey, of a supernova of mag ∼ 17 in NGC 930 (α = 2^{h}25^{m}.1, δ = +20^{o}06′, equinox 1950.0). SN 1992bf is located 9′′.4 east and 10′′.8 south of the galaxy’s center, and is somewhat embedded in the galaxy nucleus. W. Sargent and T. Small obtained spectra of this object using the 5-m Hale Telescope (+ double spectrograph) on Oct. 20 UT, showing this to be a type-I supernova.

PG 0824+289 B. Margon and P. J. Green, University of Washington; A. Klemola,

Lick Observatory; and J. G. Hoessel, University of Wisconsin, communicate: “PG 0824+289 is unique as the only known dwarf carbon (dC) star to show a composite spectrum, with a very hot DA white dwarf companion also explicitly visible in the spectrum (Heber et al., A.Ap., in press). Here we report that the object is resolved into a close pair. Elongated images are evident on the Palomar Sky Survey and archival Lick astrograph plate material. CCD imagery with the Kitt Peak 2.1-m reflector and Gunn gri filters on Oct. 17 UT reveals two objects of similar i magnitude, but quite different color, separated by 3′′.3 at position angle 120^{o}. The easternmost object is clearly the cooler one. Although spectra of the individual components are desirable, it seems reasonable to infer that the east and west resolved objects are the dC and DA, respectively. The very wide physical spacing then implied greatly complicates the evolutionary history of the pair, indicating that any recent interaction of the two stars is implausible. As many spectrograph slits are by default oriented east-west, accidentally quite close to the alignment of the two components of this system, considerable care will be needed to avoid confusion in future observations.”

NOVA SAGITTARII 1992 No. 3 Photometry by A. C. Gilmore, Mount John University Observatory

(± 0.03 unless otherwise noted): Oct. 15.42 UT, V = 7.49 ± 0.02, B–V = +0.60 ± 0.02, U–B = –0.09 ± 0.08, V–R = +0.41, V–I = +0.82; Oct. 16.40, 7.67, +0.78, +0.34 ± 0.05, +0.44, +0.87; Oct. 17.40, 7.87 ± 0.04, +0.59 ± 0.04, –0.26 ± 0.05, +0.49, +1.06.

1992 October 22 Daniel W. E. Green